OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL, PLS HELP. What is the 9th term of the geometric sequence 4, -20, 100, ...?

OpenStudy (anonymous):

A. -312,500 B. -12,500 C. 62,500 D. 1,562,500

OpenStudy (turingtest):

what is the pattern to get from one term to the next?

OpenStudy (anonymous):

-5

OpenStudy (turingtest):

great, so the formula for the nth term in the sequence \(a_n\) would be what?

OpenStudy (secret-ninja):

Just multiply the number -5 9 times. :p

OpenStudy (secret-ninja):

4 x -5 = -20 -20 x -5 = 100 100 x -5 = -500 -500 x -5 = 2500 2500 x -5 = -12500 -12500 x -5 = 62500 62500 x -5 = -312500 -312500 x -5 = 1562500

OpenStudy (secret-ninja):

There :)

OpenStudy (secret-ninja):

Do you understand?

OpenStudy (turingtest):

or you can figure out that the formula for the nth term is\[a_n=4(-5)^{n-1}\]and just plug in for n=the term you want

OpenStudy (anonymous):

That's not one of the answers so no.

OpenStudy (anonymous):

Oh never mind, yes it is

OpenStudy (secret-ninja):

Yep, @TuringTest used the easy way. :)

OpenStudy (secret-ninja):

Formulas are faster, but I take time :p lol

OpenStudy (anonymous):

Thank you so much!